Circular No. 6560

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) J. Lecacheux, Observatoire de Paris-Meudon; L. Jorda, Max-Planck-

Institut für Aeronomie; and F. Colas, Bureau des Longitudes, write: “CCD images of comet C/1995 O1, obtained with the 1.05-m telescope at Pic du Midi Observatory under good-seeing conditions on twelve nights between Jan. 12 and Feb. 10, show a dust jet of maximum length ∼ 8000 km (6′′); the length and p.a. of the jet change within 1.75 hr of observation and from one night to another, revealing the nuclear spin — the extreme position angles of 170^{o} and 230^{o} being observed on Feb. 4.28 and 10.23 UT. We infer a possible rotation period of 11.47 ± 0.05 hr from the comparison of images obtained during different nights. This jet produces a fan of aperture 70^{o} in the southwestern quadrant with concentric dust shells (first reported by O. Lardiere at Haute Provence Observatory). These shells are separated by 12 000 km (9′′) and a preliminary analysis show that they expand at ∼ 0.3 km/s (0′′.75/hr), in agreement with a spin period of ∼ 12 hr. The aperture of the fan and the appearance of the jet suggest that the source is at a latitude of −60^{o}, probably illuminated at all times, the south pole being in view at p.a. 200^{o}. The remaining linear structures might be the boundaries of fans, as suggested by Sekanina and Boehnhardt (IAUC 6542).”

Naked-eye total-magnitude and tail-length estimates: Feb. 6.22 UT, 1.8, 2^{o} (O. Skilbrei, Hoenefoss, Norway); 9.27, 1.8, 4^{o}.5 (A. Pereira, Cabo da Roca, Portugal); 11.21, 1.5, – (R. J. Bouma, Groningen, The Netherlands); 13.21, 1.3, 15^{o} (P. Candy, Viterbo, Italy); 16.22, 1.2, 5^{o} (J. Carvajal, Teruel, Spain); 17.48, 1.1, 6^{o} (G. Kronk, Troy, IL).

SUPERNOVA 1997Z IN NGC 3261 F. Winkler, A. Clocchiatti, and R. Covarrubias, Cerro Tololo Inter-

american Observatory, report a precise position for SN 1997Z from an image obtained with the 0.9-m telescope: α = 10^{h}29^{m}06^{s}.65, δ = −44^{o}39′35′′.2 (equinox 2000.0; rms 1′′), which is 62′′.7 west and 11′′.3 north of a nearby star of mag V = 14.3.

NO X-RAY TRANSIENT NEAR SMC X-1 G. Clark, Massachusetts Institute of Technology, reports: “The tran-

sient x-ray pulsar near SMC X-1, reported on IAUC 6444, does not exist. It was an artifact caused by a processing error in the received ROSAT data and is not present in the reprocessed data now in the HEASARC archive.”

1997 February 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green